# Charge ordering and spontaneous topological Hall effect in bilayer skyrmion crystals

### v1

@@ Appendix B Susceptibilities
The bare susceptibility, $\chi_{0,ij}(\mathbf{Q},i\Omega)=\expectationvalue{n_{\mathbf{Q},i}(i\Omega)n_{\mathbf{Q},j}^{\dagger}(i\Omega)}$ is equivalent to the usual density-response function. Diagrammatically, it corresponds to a generalized bubble diagram. For spin-rotation symmetric systems, the connected piece of the diagram looks like
@@
Appendix C Charge Structure Factors
In this section we elaborate on the different charge structure factors $n({{\bf{Q}}}),\Delta n({{\bf{Q}}})$ in detail, specifically in the two symmetry broken states that we find in the main text.
@@ Appendix C Charge Structure Factors
In contrast, the intra-layer charge ordered state emerging from the dispersive bands at $J_{h}=2$ shows drastic differences as plotted in Fig. 14. Here we find that the connected component of the total charge $|n({{\bf{Q}}})|$ has a sizable modulation ($\approx 0.1$) at the ${{\bf{K}}}$ points in the Brillouin zone, shown in $(b)$. Looking at the connected piece of the polarization structure factor, $\Delta n({{\bf{Q}}})$ (both real (c) and imaginary (d)), we also find peaks at the same high-symmetry points. Lastly, the imaginary part clearly displays symmetry breaking of $C_{6z}\rightarrow C_{3z}$.
Figure 13: The mean-field charge distribution in momentum space in the presence of the skyrmion with $J_{h}/t=4\,,U/t=2$, and $V/t=0$ at a filling of $\bar{n}=25/48$. (a) The absolute value of the total charge density $|n({{\bf{Q}}})|$ in the triangular Brillouin zone. (b) The connected piece corresponding to part (a). (c) The absolute value of the polarization, $|\Delta n({{\bf{Q}}})|$, and (d) the disconnected piece corresponding to (c). We see a net polarization between layers of $\approx 0.04$, with a tiny modulation with strength $\approx 0.001$ at some incommensurate momenta ${{\bf{Q}}}\neq 0$. The total charge density however looks very similar to that of the bare SkX model as shown in Fig. 5

### v2

@@ Appendix B Susceptibilities
The bare susceptibility, $\chi_{0,ij}(\mathbf{Q},i\Omega)=\expectationvalue{n_{\mathbf{Q},i}(i\Omega)n_{\mathbf{Q},j}^{\dagger}(i\Omega)}$ is equivalent to the usual density-response function. Diagrammatically, this corresponds to a generalized bubble diagram. For spin-rotation symmetric systems, the connected piece of the diagram is expressed as
@@
Appendix C Charge Orders
In this section we elaborate on the different charge orders $n({{\bf{Q}}}),\Delta n({{\bf{Q}}})$ in detail, specifically in the two symmetry broken states that we find in the main text.
@@ Appendix C Charge Orders
In contrast, the intra-layer charge ordered state emerging from the dispersive bands at $J_{h}=2$ shows drastic differences as plotted in Fig. C. Here we find that the finite $q$ component of the total charge $|n({{\bf{Q}}})|$ has a sizable modulation ($\approx 0.1$) at the $\bf{K}$}} points in the Brillouin zone, shown in $(b)$. Looking at the modulated piece of the polarization, $\Delta n({{\bf{Q}}})$ (both real (c) and imaginary (d)), we also find peaks at the same high-symmetry points. Lastly, the imaginary part clearly displays symmetry breaking of $C_{6z}\rightarrow C_{3z}$.
Figure 13: The mean-field charge distribution in momentum space in the presence of the skyrmion with $J_{h}/t=4\,,U/t=2$, and $V/t=0$ at a filling of $\bar{n}=25/48$. (a) The absolute value of the total charge density $|n({{\bf{Q}}})|$ in the triangular Brillouin zone. (b) The finite $q$ piece corresponding to part (a). (c) The absolute value of the polarization, $|\Delta n({{\bf{Q}}})|$, and (d) the uniform piece corresponding to (c). We see a net polarization between layers of $\approx 0.04$, with a tiny modulation with strength $\approx 0.001$ at some incommensurate momenta ${{\bf{Q}}}\neq 0$. The total charge density however looks very similar to that of the bare SkX model as shown in Fig. 5